OpenStudy (thadds2003):

Medal and Fan. :/ Combine like terms. 8 + 6x – 11x + 12 A. 20 – 17x B. 14x + 1x C. 20 + 5x D. 20 – 5x

pooja195 (pooja195):

8 + 6x – 11x + 12 x=pooja's numbers without variables=thadds combine the Pooja's then combine thadds put em together and u get your answer

OpenStudy (thadds2003):

wait what?

pooja195 (pooja195):

Ok x's are pooja so let's put pooja with pooja :) 6x-11x=?

OpenStudy (thadds2003):

i dont get it, -5?

pooja195 (pooja195):

Yeah -5x combine the numbers without the variables 8+12=?

OpenStudy (thadds2003):

20

pooja195 (pooja195):

Good so we have -5x+20

pooja195 (pooja195):

Or the flip 20-5x

OpenStudy (thadds2003):

Ok thx here medal and fan.

pooja195 (pooja195):

;p thx